please explain

How many sublevels are in the n 3 evel? Number sub
Answer #1

1)

if n = 3

number of sublevels l = n-1 = 3-1 = 2

2)

orbitals in n = 3

subshells = 3s, 3p and 3d

3s = 1 orbital 3p = 3 orbitals 3d = 5 orbitals = 1+3+5 = 9 orbitals!

3)

Max electrons in n = 3 level

3s = 2 electrons

3p = 6 electrons

3d = 10 electrons

total electronws 2+6+10 = 18 electrons